Could you please tell me whether or not the equations are linear and why?
x=y, 6=x , xy=5

Answer by Theo(13342)

If you can convert it to a form of:

ax + by = c

or:

y = mx + b

then it's linear.

x = y is in the form of y = mx + b.

6 = x is already in the form of ax + by = c

It's the same as x = 6.

xy = 5 is probably not linear.

you get y = 5/x which doesn't conform to a linear equation form.

graphs of these equations can be seen by clicking on the following hyperlink.

Picture of graph of equations

y = 5/x is definitely not linear.